40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15717     Entry time: Wed Dec 9 11:54:11 2020
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ITMX HeNe replaced 

The ITMX Oplev (installed in March 2019) was near end of life judging by the SUM channel (see Attachment #1). I replaced it yesterday evening with a new HeNe head. Output power was ~3.25 mW. The head was labelled appropriately and the Oplev spot was recentered on its QPD. The lifetime of ~20 months is short but recall that this HeNe had already been employed as a fiber illuminator at EX and so maybe this is okay.

Loop UGFs and stability margins seem acceptable to me, see Attachment #2-#3.

Attachment 1: OLtrend_old_ndscope.png  200 kB  | Hide | Hide all
OLtrend_old_ndscope.png
Attachment 2: ITMX_OL_P.pdf  17 kB  | Hide | Hide all
ITMX_OL_P.pdf
Attachment 3: ITMX_OL_Y.pdf  18 kB  | Hide | Hide all
ITMX_OL_Y.pdf
ELOG V3.1.3-